# 1bit-Merging: Dynamic Quantized Merging for Large Language Models

### v1

@@ Abstract
Recent advances in large language models have led to specialized models excelling in specific domains, creating a need for efficient model merging techniques. While traditional merging approaches combine parameters into a single static model, they often compromise task-specific performance. However, task-specific routing methods maintain accuracy but introduce substantial storage overhead. We present 1bit-Merging, a novel framework that integrates task-specific routing with 1-bit quantized task vectors to balance performance and storage efficiency. Our approach leverages the observation that different task-specific models store knowledge in distinct layers—chat models primarily in attention layers and math/code models in MLP layers—enabling targeted compression strategies. Through extensive experiments with LLaMA2 and Mistral model families across chat, mathematical reasoning, and code generation tasks, we demonstrate that 1bit-Merging achieves comparable or superior performance to existing methods while significantly reducing storage requirements. Our framework offers a practical solution for combining specialized models while maintaining their individual strengths and addressing the storage challenges of current approaches.
@@
Shuqi Liu1,2, Han Wu2,, Bowei He1, Zehua Liu2, Xiongwei Han2, Mingxuan Yuan2, Linqi Song1, 1 Department of Computer Science, City University of Hong Kong 2 Huawei Noah’s Ark Lab {shuqiliu4-c, boweihe2-c}@my.cityu.edu.hk wu.han1@huawei.com linqi.song@cityu.edu.hk
@@ 1 Introduction
Large language models have achieved remarkable progress, demonstrating strong performance on a wide range of tasks Touvron et al. (2023); Zhao et al. (2023). As researchers continue to fine-tune these models for specific domains, there is a growing need to combine their specialized capabilities into a single model (Yang et al., 2024; Goddard et al., 2024). While multi-task learning offers one solution Sanh et al. (2022); Fifty et al. (2021), it requires extensive computational resources and simultaneous access to all task-specific datasets. Recent advances in parameter-space model merging Wortsman et al. (2022); Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024b) provide an efficient alternative - by directly operating on model parameters, these methods preserve data privacy and eliminate the need for expensive retraining.
Traditional model merging approaches Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024b) typically combine the parameters of multiple fine-tuned models, or expert models, into a single static model without additional training, thereby enabling efficient multi-task functionality. However, merging models from different domains often sacrifices task-specific performance, resulting in a noticeable gap compared to individual expert models. In contrast, merging with task-specific routing (Muqeeth et al., 2024; Lu et al., 2024) dynamically prioritizes relevant task vectors based on input data, effectively maintaining accuracy by isolating task-specific parameters. However, this routing-based merging strategy introduces substantial storage overhead, as it necessitates the preservation of all task vectors to ensure task relevance and performance. Thus, despite their ability to uphold model accuracy, task-specific routing methods face severe storage challenges, limiting their scalability and practicality in resource-constrained environments.
@@
To effectively balance performance and storage efficiency, we introduce 1bit-Merging, a novel dynamic merging framework that integrates task-specific routing with 1-bit quantized task vectors. Recognizing the substantial redundancy inherent within task vectors, we implement 1-bit compression, which significantly reduces storage requirements without notably compromising the model’s effectiveness. Notably, we observe that different task-specific models store knowledge in distinct layers. For chat-oriented models, knowledge is predominantly stored in attention layers, enabling us to compress MLP layers. Conversely, for math and code-related models, knowledge is primarily stored in MLP layers, allowing us to compress attention layers. Building upon the compressed task vectors, our framework employs task-specific routing to establish a task-specific base model. This base model serves as the foundation for integrating the remaining compressed task vectors, ensuring that each task leverages the most relevant and efficient parameters. 1bit-Merging thus offers a balanced solution that maintains the performance advantages of task-specific routing while addressing the storage inefficiencies of existing approaches.
@@
To empirically demonstrate the effectiveness of 1bit-Merging, we conduct extensive experiments by combining it with existing model merging approaches. We merged three widely adopted fine-tuned models—specializing in general knowledge (Chat), mathematical reasoning (Math), and code generation (Code)—derived from the LLaMA2-7B/13B and Mistral 7B families. Through extensive experiments across multiple tasks, we demonstrate that our approach not only outperforms traditional model merging methods but also achieves better storage efficiency than task-specific routing approaches. Our task vector compression method effectively preserves and often enhances the capabilities of fine-tuned models. By integrating these compressed task vectors with our dynamic routing strategy, our 1bit-Merging method achieves superior performance. As illustrated in Figure 1, compared to individually fine-tuned models, our 1bit-Merging method delivers better performance across all domains.
@@
To sum up, our contributions include: (1) We propose a novel dynamic merging framework that integrates task-specific routing with 1-bit quantized task vectors. (2) We empirically demonstrate that different task-specific models store knowledge in distinct layers, enabling targeted compression strategies based on specific compression position. (3) Through comprehensive evaluations, we validate that our proposed method enhances model merging performance across various domains and achieves better storage efficiency than task-specific routing approaches.
@@ 2 Related Work
Modeling merging (Yang et al., 2024; Goddard et al., 2024) has attracted much attention for its several advantages: 1) it significantly reduces the storage and deployment expenses by consolidating multiple models into a single one (Ilharco et al., 2023; Yu et al., 2024a); 2) it operates in a plug-and-play manner, eliminating the need for additional training(Wortsman et al., 2022; Yu et al., 2024b); and 3) it can improve performance on individual tasks(Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023) while also improving out-of-domain generalization(Cha et al., 2021; Jin et al., 2023). Based on whether the merged model remains consistent across all samples or tasks, model merging approaches can be categorized into two types: static model merging and dynamic model merging.
2.1 Static Model Merging
Static model merging primarily explores general strategies for combining models, such as Average Merging (Wortsman et al., 2022), Task Arithmetic (Ilharco et al., 2023), and Ties-Merging (Yadav et al., 2023). Notably, Wortsman et al. (2022) first demonstrated that even a straightforward weight averaging of base models can improve both the performance and robustness of downstream tasks. Building on this, Task Arithmetic (Ilharco et al., 2023) refines the merging process by introducing task vectors, proposing that simple arithmetic operations on these vectors can effectively modify models and yield a better merged model. Expanding upon the concept of task vectors, methods like DARE (Yu et al., 2024b) and Ties (Yadav et al., 2023) adopt pruning-then-scaling techniques to merge task vectors, based on the premise that not all parameters equally contribute to the final performance. However, static merging of models from different domains often sacrifices task-specific performance to strike a balance between generalization capacity and task-specific effectiveness. To this end, we explore the dynamic model merging in this work.
2.2 Dynamic Model Merging
@@
Table 1: Impact of different layer types on compression performance across Math, Code, and Chat expert models derived from LLaMA-2 7B model. Attention layer compression performs best for Math and Code models, while MLP layer compression yields optimal results for Chat model.
3 Method
3.1 Traditional Model Merging and Merging with Task-Specific Routing
We first compare traditional model merging with the merging strategies combined with task-specific routing. Starting with $K$ fine-tuned models $\{\theta^{t_{1}}_{\text{SFT}},\theta^{t_{2}}_{\text{SFT}},\ldots,\theta^{t_{K}}_{\text{SFT}}\}$ derived from a common pre-trained backbone $\theta_{\text{PRE}}$, each task vector is defined as the difference between and after finetuning:
$\delta_{t_{k}}=\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}},\quad\text{for }k\in\{1,\ldots,K\}.$
Traditional merging aggregates all task vectors to construct a single static merged model:
$\theta_{\text{merged}}^{\text{traditional}}=\theta_{\text{PRE}}+\sum_{k=1}^{K}\delta_{t_{k}}.$
In contrast, merging with routing leverages task specialized knowledge tailored to each input $\mathbf{x}$ through a router:
$k^{*}=\text{Router}(\mathbf{x}),\quad\theta_{\text{merged}}^{\text{routing}}=\theta_{\text{PRE}}+\delta_{t_{k^{*}}}.$
As shown in Figure 2(a), traditional merging methods like Task Arithmetic Ilharco et al. (2023) are highly impacted by task interference: merging dissimilar tasks (math and chat) degrades performance more than similar tasks (math and code), and adding more tasks exacerbates the decline. In contrast, task-specific routing prevents interference and preserves the accuracy of fine-tuned models.
(a) Gaps from merging with task-specific routing.
(b) Impact of merging with the fine-tuned model as base.
Figure 2: Performance of merged model in traditional model merging and merging with task-specific settings.
3.2 Impact of Base Model in Model Merging
Building on the strength of task-specific routing, we further investigate the impact of selecting different base models in model merging. Specifically, we compare the performance outcomes when employing a pre-trained backbone $\theta_{\text{PRE}}$ versus a math-finetuned model $\theta_{\text{Math}}=\theta_{\text{PRE}}+\delta_{t_{\text{Math}}}$ as the base model for mathematical input data. As shown in Figure 2(b), substituting the pre-trained backbone with a math-finetuned model significantly enhances performance on the GSM8K Cobbe et al. (2021) dataset. Consequently, initializing the base model with task-specific fine-tuning allows the merged model to more effectively solve the corresponding task. Despite surpassing individually fine-tuned models, utilizing a task-specific base model concurrently increases storage requirements due to the necessity of loading multiple task vectors.
@@ 3.2 Impact of Base Model in Model Merging
To balance performance with storage efficiency, we introduce 1bit-Merging, a dynamic merging method that combines task-specific routing with 1-bit quantized task vectors.
Compression of Task Vectors.
Recognizing the substantial redundancy within task vectors, we apply 1-bit weight quantization to the task vectors, converting the weight matrices in Linear layers from FP32/16 precision to a 1-bit format. In this quantization process, each element of the task vector is set to either +1 or -1. To maintain performance despite the aggressive compression, we scale the binary weight matrices with scalar values in FP16 format. This scaling ensures that the quantized weights preserve the original weight $L_{2}$ norm, thereby maintaining model performance after the extremely low-bit compression Liu et al. (2024). The scaling factor $\alpha$ is computed as:
$\alpha=\frac{\|\mathbf{W}\|_{1}}{{m\cdot n}}$
@@
Compression Position.
We examine 1-bit compression effectiveness across different model components by selectively compressing Attention layers, MLP layers, and all Linear layers. As shown in Table 1, each model type exhibits distinct compression characteristics. Math and Code models achieve optimal performance when only attention layers are compressed while Chat models perform best with MLP layer compression. Remarkably, the compressed versions of both Math and Code models demonstrate slight performance improvements over their original fine-tuned counterparts These patterns reveal a fundamental difference in how knowledge is distributed within the model architecture: task-specific capabilities (like mathematical reasoning and code generation) primarily reside in MLP layers, while attention layers appear to store more general, transferable knowledge. Based on these insights, we strategically apply attention layer quantization to Math and Code models, and MLP layer quantization to Chat models when compressing task vectors.
Dynamic Routing and Merging.
Our merging framework then incorporates the compressed task vectors with a task-specific routing mechanism, where a trained router analyzes the input data $\mathbf{x}$ to produce a probability distribution $\mathbf{p}$ across different tasks. The most relevant task vector $\delta_{t_{k}^{*}}$ is selected based on the highest probability $k^{*}$ and added to the pre-trained model parameters $\theta_{\text{PRE}}$ to form a task-specific base model $\theta_{\text{base}}=\theta_{\text{PRE}}+\delta_{t_{k}^{*}}$. Finally, we apply Ties Merging Yadav et al. (2023) to integrate the remaining compressed task vectors into $\theta_{\text{base}}$, resulting in a comprehensive model that dynamically adapts to the most relevant tasks while maintaining overall performance through efficient parameter integration.
4 Experiments
4.1 Experimental Setup
Baselines.
We evaluate the effectiveness of our 1bit-Merging method through a two-step process. First, we assess the performance of compressed task vectors by comparing compressed expert models— which integrate these vectors with the pretrained backbone—against the individually fine-tuned models. Second, we benchmark 1bit-Merging against existing model merging methods: Task Arithmetic, Ties-Merging, and DARE. Task Arithmetic (Ilharco et al., 2023) enhances the merging process by introducing task vectors, suggesting that simple arithmetic operations on these vectors can effectively merge models. Building on the concept of task vectors, both DARE (Yu et al., 2024b) and Ties-Merging (Yadav et al., 2023) employ pruning-then-scaling methods to merge task vectors, based on the assumption that not all parameters contribute equally to the final performance.
@@ 4.1 Experimental Setup
Table 2: Performance evaluation of merged LLaMA2-7B Models (Chat, Math, Code) across 7 task-specific datasets

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have led to specialized models excelling in specific domains, creating a need for efficient model merging techniques. While traditional merging approaches combine parameters into a single static model, they often compromise task-specific performance. However, task-specific routing methods maintain accuracy but introduce substantial storage overhead. We present 1bit-Merging, a novel framework that integrates task-specific routing with 1-bit quantized task vectors to balance performance and storage efficiency. Our approach further leverages the observation that different task-specific models store knowledge in distinct layers, such as MLP and attention layers, enabling targeted compression strategies. Through extensive experiments with LLaMA2 and Mistral model families across general knowledge, mathematical reasoning, and code generation tasks, we demonstrate that 1bit-Merging not only outperforms traditional model merging methods but also achieves better storage efficiency than task-specific routing approaches, maintaining 94.53% of their performance while reducing storage requirements to 55.02% on merging Chat, Math and Code fine-tuned models on Mistral 7B models.
@@
Shuqi Liu1,2, Yuxuan Yao1, Bowei He1, Zehua Liu2, Xiongwei Han2, Mingxuan Yuan2 Han Wu2,, Linqi Song1, 1 Department of Computer Science, City University of Hong Kong 2 Huawei Noah’s Ark Lab shuqiliu4-c@my.cityu.edu.hk wu.han1@huawei.com linqi.song@cityu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress, demonstrating favourable performance on a wide range of tasks Touvron et al. (2023); Zhao et al. (2023). As a variety of fine-tuned (FT) models for specific domains have emerged, there is a growing need to combine their specialized capabilities into a single model (Yang et al., 2024; Goddard et al., 2024). While multi-task learning offers one solution Sanh et al. (2022); Fifty et al. (2021), it requires extensive computational resources and simultaneous access to all task-specific datasets. Recent advances in parameter-space model merging Wortsman et al. (2022); Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024) provide an efficient alternative, which combines pre-trained (PT) and FT models through parameter manipulation without additional training overhead, resulting in a unified model with comprehensive capabilities.
Within the framework of model merging, task vectors Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024) are pivotal for encoding task-specific knowledge. Defined as parametric deltas between PT and FT model weights, these vectors theoretically facilitate cross-model capability integration via linear arithmetic operations. Nonetheless, traditional merging often sacrifices task-specific performance, resulting in a noticeable gap compared to individual expert models. For instance, merging the math-specific, and code-specific models may result in a model that performs better in math than the code-specific model, but worse in code tasks. To alleviate such challenges, merging with task-specific routing (Muqeeth et al., 2024; Lu et al., 2024) dynamically prioritizes relevant task vectors based on input data, effectively maintaining accuracy by isolating task-specific parameters. However, this routing-based merging strategy introduces substantial storage overhead, as it necessitates the preservation of all task vectors. SVD-based and quantization-based approaches are proposed to resolve the problem Wang et al. (2025); Liu et al. (2024), but require complex high-dimensional computations or utilize the same quantization method for all components, contradicting previous observations Xu et al. (2025); Liu et al. (2025) that different layers possess varying importance. Building upon earlier studies that highlight the varying sensitivity across layers (Liu et al., 2025), we further investigate and identify differing levels of importance among functional modules, such as MLP and Attention, which can be leveraged for model merging. Notably, our empirical findings reveal that task-specific models encode knowledge in distinct modules. For example, chat-oriented models predominantly store knowledge within attention layers, enabling the compression of MLP layers. In contrast, math- and code-related models primarily encode knowledge in MLP layers, allowing for the compression of attention layers.
@@
In this work, we introduce 1bit-Merging, a novel dynamic merging framework that integrates task-specific routing with adaptive 1-bit quantized task vectors. Recognizing the substantial redundancy inherent within task vectors, we implement 1-bit compression, which significantly reduces storage requirements without notably compromising model capacities. Building upon the compressed task vectors, our method employs task-specific routing to establish a task-specific base model. This base model serves as the foundation for integrating the remaining compressed task vectors, ensuring that each task leverages the most relevant and efficient parameters. 1bit-Merging thus offers a balanced solution that maintains the performance advantages of task-specific routing while addressing the storage inefficiencies in existing methods.
@@
Empirically, we conducted extensive experiments by comparing 1bit-Merging with existing model merging approaches. We merged three widely used fine-tuned models focused on general knowledge (Chat), mathematical reasoning (Math), and code generation (Code) from the LLaMA2 and Mistral families. Our results show that 1bit-Merging outperforms traditional model merging methods while offering better storage efficiency than task-specific routing approaches. The task vector compression method preserves and often enhances the capabilities of FT models. By combining these compressed task vectors with our dynamic routing strategy, 1bit-Merging achieves superior performance. As shown in Figure 1, it surpasses individually fine-tuned models across all domains.
@@
To sum up, our contributions include: (1) We discover that different task-specific models store knowledge in distinct layers, enabling targeted compression strategies based on specific compression position. (2) We propose a novel dynamic merging framework that integrates task-specific routing with 1-bit quantized task vectors. (3) Through comprehensive evaluations, we demonstrate that 1bit-Merging not only outperforms conventional merging methods but also achieves superior storage efficiency compared to task-specific routing. It retains 94.53% of the performance of full-precision models while reducing storage requirements to 55.02%, when merging on Mistral 7B models.
@@ 2 Related Work
Model merging, a paradigm to enhance model generalization and robustness by synthesizing heterogeneous neural networks, is broadly categorized into static and dynamic approaches.
Static Model Merging
Static model merging primarily explores general strategies for combining models, such as Average Merging (Wortsman et al., 2022), Task Arithmetic (Ilharco et al., 2023), and TIES-Merging (Yadav et al., 2023). Notably, Wortsman et al. (2022) first demonstrated that even a straightforward weight averaging of base models can improve both the performance and robustness of downstream tasks. Building on this, Task Arithmetic (Ilharco et al., 2023) refines the merging process by introducing task vectors, proposing that simple arithmetic operations on these vectors can effectively modify models and yield a better merged model. Expanding upon the concept of task vectors, methods like DARE (Yu et al., 2024) and TIES-Merging (Yadav et al., 2023) adopt pruning-then-scaling techniques to merge task vectors, based on the premise that not all parameters equally contribute to the final performance. However, static merging of models from different domains often sacrifices task-specific performance to strike a balance between generalization capacity and task-specific effectiveness. To this end, we explore the dynamic model merging in this work.
Dynamic Model Merging
@@
3 Method
3.1 Preliminaries
Task Vectors
Given a set of $K$ fine-tuned models $\{\theta^{t_{1}}_{\text{SFT}},\theta^{t_{2}}_{\text{SFT}},\ldots,\theta^{t_{K}}_{\text{SFT}}\}$ derived from a shared pre-trained backbone $\theta_{\text{PRE}}$, each task vector is defined as the parameter difference between the fine-tuned and pre-trained models:
$\Delta\theta^{t_{k}}=\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}}.$ (1)
Traditional merging methods combine all task vectors to produce a single, static merged model:
$\theta_{\text{merged}}^{\text{traditional}}=\theta_{\text{PRE}}+\sum_{k=1}^{K}\delta_{t_{k}}.$ (2)
In contrast, routing-based merging dynamically selects a task-specific vector based on the input $\mathbf{x}$ using a router function:
$\theta_{\text{merged}}^{\text{routing}}=\theta_{\text{PRE}}+\delta_{t_{i}},\quad i=\mathop{\mathrm{argmax}}\limits_{k}~{}\mathrm{Router}(\mathbf{x}),$ (3)
where $t_{i}$ denotes the task vector selected by the well-trained router.
Figure 2: Performance comparison of different strategies on GSM8K on LLaMA2-7B series. TIES-Merging achieves superior performance when choosing Math FT model as base model.
Impact of Base Model
Leveraging the advantages of task-specific routing, we further investigate how the choice of base model affects model merging performance. Specifically, we compare the use of a pre-trained (PT) backbone $\theta_{\text{PRE}}$ versus a fine-tuned (FT) model for mathematics $\theta_{\text{Math}}=\theta_{\text{PRE}}+\delta_{t_{\text{Math}}}$, as the base model in mathematical tasks. As shown in Figure 2, substituting the PT backbone with a math-specialized FT model (Math-FT) yields substantial performance improvements on the GSM8K dataset Cobbe et al. (2021) when used in conjunction with the TIES-Merging approach. This result highlights the importance of selecting an appropriate base model for effective model merging.
Despite outperforming individually fine-tuned models, the use of a task-specific base model introduces increased storage overhead, as multiple task vectors requires to be loaded concurrently.
3.2 Unveiling the Varying Module Importance
Inspired by prior work on module-level and layer-level importance in LLMs Meng et al. (2022); Liu et al. (2025), we investigate how different expert models respond to 1-bit quantization across various architectural components, including the Attention module, MLP module, and all Linear layers.
As shown in Table 1, for Chat model, performance remains largely preserved after quantizing the MLP module, suggesting that instruction-following capabilities are primarily encoded in the Attention module. In contrast, for Math model, quantizing the Attention module has minimal impact, while quantizing MLP leads to significant degradation—indicating that complex reasoning capabilities are predominantly stored in MLP subnetworks. Similar phenomenon can be observed in Table 3 and Table 11 for Mistral-7B and LLaMA2-13B models correspondingly. These findings highlight that knowledge localization varies across modules in a task-dependent manner, and that applying a task-agnostic compression strategy—such as uniform quantization—can severely degrade performance on complex reasoning tasks.
@@ 3.2 Unveiling the Varying Module Importance
Table 1: Impact of quantizing different module types on the performance of compressed models across Math, Code, and Chat expert variants derived from the LLaMA-2 7B backbone. Compression of attention modules yields the best results for Math and Code models, while MLP module compression performs optimally for the Chat model.
@@
Based on the aforementioned analysis, we introduce 1bit-Merging to balance performance with storage efficiency, which is a dynamic merging framework that combines task-specific routing with 1-bit quantized task vectors. According to the discussion in Section 3.2, we only quantify specific modules, the workflow is outlined as follows.
1-bit Quantization of Task Vectors
Recognizing the substantial redundancy within task vectors, we apply 1-bit weight quantization to the task vectors, converting the weight matrices in specific modules from FP32/16 precision to a 1-bit format. In this quantization process, each element of the task vector is set to either +1 or -1. To maintain performance despite the aggressive compression, we scale the binary weight matrices with scalar values in FP16 format. This scaling ensures that the quantized weights preserve the original weight $L_{2}$ norm, thereby maintaining model performance after the extremely low-bit compression Liu et al. (2024). The scaling factor $\alpha$ is computed as:
$\alpha=\frac{\|\mathbf{W}\|_{1}}{{m\cdot n}},$ (4)
@@
Dynamic Routing and Merging
Our merging framework dynamically selects and integrates compressed task vectors through a task-aware routing mechanism. Given an input $\mathbf{x}$, a lightweight router—implemented as a three-layer linear network with Leaky ReLU activations and batch normalization—produces a softmax-normalized probability distribution over tasks:
$\mathbf{p}=\text{softmax}(f_{\theta}(\mathbf{x})),$ (6)
where $f_{\theta}$ denotes the router’s neural architecture. The dominant task category is selected via:
$k^{*}=\mathrm{argmax}(\mathbf{p}),$ (7)
identifying the most relevant task vector $\tilde{\delta}_{t_{k^{*}}}$. This vector is then added to the pre-trained backbone $\theta_{\text{PRE}}$ to construct a task-specific base model:
$\theta_{\text{base}}=\theta_{\text{PRE}}+\tilde{\delta}_{t_{k^{*}}}.$ (8)
Subsequently, we apply TIES-Merging Yadav et al. (2023) to further integrate the remaining compressed task vectors into $\theta_{\text{base}}$, yielding a final merged model that adapts dynamically to the input while preserving overall performance through efficient parameter fusion:
$\mathbf{M}_{\text{merged}}=\text{TIES}\left(\theta_{\text{base}},\{\delta_{t_{k}}\mid k\neq k^{*}\}_{k=1}^{K}\right).$ (9)
We use TIES-Merging to integrate task vectors, as it effectively resolves parameter conflicts via sign-based consensus, which aligns well with our use of 1-bit quantized task vectors.
@@
Table 2: Performance evaluation of merged LLaMA2-7B Models (Chat, Math, Code) across 7 task-specific datasets.
4 Experiments
In this section, we present a comprehensive evaluation of our 1bit-Merging method from multiple perspectives. We begin by comparing our approach to expert models and traditional model merging techniques, aiming to assess its ability to preserve and effectively combine knowledge across different tasks. We then evaluate its performance against dynamic merging methods, focusing on adaptive and task-aware fusion scenarios. Finally, we examine the scalability of our method with respect to model architecture and parameter size, demonstrating its broader applicability across diverse models.
4.1 Experimental Setup
Baselines
We consider both traditional and dynamic model merging methods as baselines. For traditional approaches, we include Task Arithmetic (Ilharco et al., 2023), which linearly combines task-specific models in the weight space; TIES-Merging Yadav et al. (2023), which identifies consensus directions for parameter fusion; and DARE Yu et al. (2024), which randomly prunes parameters before merging to improve generalization. As a representative dynamic merging method, we adopt Twin-Merging Lu et al. (2024), which introduces a trainable router to adaptively combine models based on input features. In particular, Twin Merging employs singular value decomposition (SVD) to compress the parameter space, enabling efficient and expressive model interpolation. We refer readers to Appendix A.1 for detailed implementation of our router, and hyperparameter settings for the merging methods.
General Knowledge Mathmetical Reasoning Code Generation
/ 59.05 65.97 55.69 42.53 9.16 49.6 42.7 46.37
Attention 53.25 57.03 54.59 39.20 8.14 46.6 36.0 42.12
MLP 54.71 58.00 53.24 38.51 8.14 47.6 35.4 42.23
Chat Linear 54.95 58.83 53.73 38.82 8.26 47.6 37.2 42.77
/ 60.77 58.68 44.68 63.38 22.74 38.1 23.8 44.59
Attention 60.11 58.48 41.13 58.76 19.92 41.1 25.6 43.59
MLP 58.74 58.12 43.57 53.37 16.68 44.9 29.3 43.53
Math Linear 57.93 58.24 39.41 47.16 13.9 41.1 25.6 40.48
/ 50.58 53.19 45.29 31.69 4.84 50.9 40.9 39.63
Attention 47.88 53.51 42.96 27.29 5.16 41.6 33.5 35.99
MLP 45.65 54.56 42.59 34.34 6.08 47.9 28.0 37.02
Code Linear 39.99 53.80 36.47 22.52 4.82 40.6 31.1 32.76
Table 3: Impact of quantizing different module types on Mistral 7B. Compression of attention modules yields the best results for Math and Code models, while Linear module compression performs optimally for the Chat model.
Models & Dataets
Our experimental evaluation focuses on three model families: LLaMA-2 7B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and LLaMA-2 13B Touvron et al. (2023), each covering distinct specializations in: general knowledge (Chat), mathematical reasoning (Math), and code generation (Code). Additional details regarding the model variants and their sources are provided in Appendix A.2.
We assess performance using seven benchmark datasets across three domains: MMLU Hendrycks et al. (2020), HellaSwag Zellers et al. (2019) and TruthfulQA Lin et al. (2022) for assessing general knowledge and reasoning capabilities; GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) for testing mathematical reasoning proficiency; HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) for evaluating code generation ability. To ensure consistent and unbiased assessment, model performance is evaluated using zero-shot accuracy, with pass@1 rate specifically measuring code generation correctness.
Model MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval Average
Chat (1bit) 46.40 58.23 43.70 22.74 5.08 0.0 0.0 25.55
Chat (SVD, $r=128$) 46.52 58.04 45.17 21.99 4.94 0.0 0.0 25.14
Math (1bit) 41.03 56.68 34.52 47.46 8.42 23.1 14.0 32.17
Math (SVD, $r=128$) 40.76 57.60 33.29 45.49 7.62 22.8 14.0 31.65
Code (1bit) 41.36 57.81 32.80 8.19 3.36 27.3 11.0 25.97
Code (SVD, $r=128$) 33.17 57.23 31.65 7.58 3.06 26.6 4.9 24.94
Table 4: Performance of compressed expert models (Chat, Math, Code) under 1-bit quantization and SVD compression (r=128) applied to task vectors for LLaMA2-7B models.
4.2 Compare to Expert Models and Traditional Merging
Using Chat, Math, and Code fine-tuned variants of the LLaMA2-7B pretrained model, we first evaluate the performance of compressed expert models against their fine-tuned counterparts to assess the effectiveness of task-vector compression. We then compare our 1bit-Merging model with existing task-vector-based merging methods. Results across seven datasets are summarized in Table 2.
Compressed Expert Models
The compressed expert models achieve strong performance on their respective specialized tasks and exhibit varying degrees of cross-domain generalization. As shown in Table 1, the 1bit-Chat model achieves an average score of 25.55, slightly outperforming its fine-tuned counterpart (25.50). The 1bit-Math model improves from 31.52 to 32.17, while the 1bit-Code model increases from 24.98 to 25.97. These results suggest that module-specific 1-bit quantization preserves—and in some cases even enhances—the overall model performance across various tasks. We think this is because the quantization eliminates the impact of insignificant features.
Moreover, we further observe that the compressed models maintain strong performance beyond their primary domain. For instance, the 1bit-Math model also improves by 1.11 points on general knowledge and 1.25 points on code generation—areas outside its main focus. This suggests that the removal of redundant domain knowledge by our module-specific compression strategy could improve the out-of-domain capacities while preserving the performance in the original domain, without introducing additional parameters or computational overhead.
Compare to Traditional Merging
As presented in Table 2, our 1bit-Merging approach achieves superior performance over traditional model merging techniques along two key dimensions. (1) Broad and balanced performance across domains: While baseline methods exhibit domain-specific strengths—TIES-Merging performs well on mathematical tasks, and DARE achieves strong results in general knowledge and code generation—our approach delivers consistently competitive performance across all domains without bias toward any single task. Specifically, 1bit-Merging outperforms TIES-Merging by 1.45 points in mathematical reasoning and DARE by 2.71 points in general knowledge. Although we observe a modest 0.75-point drop in code generation compared to DARE, this is primarily due to the relatively low performance of the fine-tuned Code model itself. In fact, our method improves upon that baseline by 8.35 points, demonstrating substantial gains in absolute terms. (2) Outperforms average expert performance: Most notably, our method achieves an average score of 36.74 across all tasks, exceeding the average performance (35.16) of individual fine-tuned expert models. This gain of 1.58 points suggests that our merging strategy effectively preserves and combines specialized capabilities from different experts, enabling overall performance that surpasses the best individual models.